OWG Meeting 20-Jun-2012
- Review of minutes from 6-Jun-2012 meeting
- Intel-based ROC purchase - Dave Abbott, Hovanes, Elliott
- how many do we need?
- Plans for Tagger area - Hovanes
- CAEN FPGA boards - Hovanes
- JInventory database and barcode scanners - Hovanes, Elliott
- EVIO status - Carl, Elliott
- Counting House status - Hovanes
- 12GeV Schedule - Hovanes, David, Elliott
- Elog and Accelerator division projects - Graham
- DAQ group - Bryan, Dave A, Graham
- Raw data tape
- Java GUI connection to cMsg
- CODA3, EVIO, ECS/CSS, raw event file, disentangler, readout list
- EPICS - Hovanes, Vanik
- Controls - Elliott
- Solenoid Controls Redesign, FastTrack schedule, PXI status, FMEA, P&I, vacuum, connection to Cryo group PLC, etc.
- Electronics, trigger and DAQ status reports - Chris, Alex, Dave, Fernando, Ben, Ed, William, Bryan
- Mantis Task Tracker
- Midas Elog
- Calibration/Alignment issues?
- Computer system manager
1:30 PM Wed 20-Jun-2012 CC F326
Note...no meeting 4-Jul
1:30 PM Wed 18-Jul-2012 CC F326
Present: Elliott W, Mark I, David L, Simon T, Bryan M, Hovanes E, Vanik K, Beni Z.
Hovanes reported we need 75 ROCs total and have 11 duals. We plan to purchase quads for the remainder assuming the price comes in as expected. Glenn and Dave A will put out an RFQ for the entire lab purchase soon.
Hovanes plans to order two CAEN general purpose FPGA boards V1495 soon. CLAS used them successfully and Elliott put three of them in the budget many years ago. They should replace a wall of NIM logic. They have 64 inputs and 32 outputs, with various options for each. We need to learn how to program them. (later...Ben R will give us a tutorial later this summer...ejw).
JInventory and barcode scanner
The sales rep loaned us a wireless scanner a day or so ago. So far it seems to work but we need to test it on the small barcode tags we need to use.
Hovanes reported the UPS and RDC are installed. Giles has completed the plans for power under the floor. The techs are ordering parts now. Installation should commence shortly.
All online and control planning lines are due. Elliott and Dave have completed all online lines other than two that are intimately bound up with the trigger. Not that the trigger lines are not due until the end of the year, so there is a mismatch. Elliot and Dave just need to complete the writeups. Hovanes is making progress on many of controls planning lines as well.
Simulated raw data tape
With a data challenge planned for the fall Elliott talked with Graham to get this back on track again, and Graham agreed. We need this by mid-summer or so.
codaLite for FDC test
Beni currently plans to use codaLite for the early stages of the FDC test. Elliott will upgrade the InterruptService classes to work with the new Linux-based ROCs and will use Bryan's library facilities.
Temperature and humidity measurements
The BCAL group is happy with Josh's proposal to read out 96 RTD and 24 temp/humidity sensor channels via Allen-Bradley CompactLogix PLC systems. We will purchase two systems, one for upstream and the other for downstream. This will leave room for expansion and for the CDC and Start Counter thermocouples. Tom will decide where the two chassis will be located.
Vanik continues to work on the EPICS driver. He needs to extend it to handle synch pulses.
Elliott is creating a Fasttrack schedule for the solenoid test. Currently cooldown is scheduled for Nov and we need to know if this is possible. If not we need to know what the critical path items are.
Bryan reported that SD's are due soon in two batches. FADC250's are due in mid-Aug. Not sure about FADC125's.